### v1

# Thin films spreading under the action of surface acoustic waves

Abstract
Experimentally and theoretically, we consider the problem of a spreading drop of millimetric thickness driven by a surface acoustic wave (SAW). The experimental results show that, after initial transients, the drop spreads with approximately constant speed while leaving a thin layer of fluid behind. To better understand the system, we develop a theoretical model that couples acoustic driving with the fluid mechanics of a spreading drop. An important feature of the problem is that the strength of the acoustic driving depends on the fluid thickness, with strong attenuation of the forcing in regions of the domain covered by fluid. The presented model provides results consistent with the experiments, particularly regarding drop height and spreading speed. The constant spreading speed observed in the experiments is also captured by an approximate traveling-wave solution, which provides predictions for the dependence of drop thickness and spreading speed on the acoustic amplitude.
1 Introduction
Dynamic wetting of a solid substrate by a film of liquid is a common occurrence. In natural systems it appears in surfactant films that wet the lungs to promote breathing and the eyes to keep them moist (Veldhuizen & Haagsman, 2000; Holly & Lemp, 1971), or in the spreading of water drops on solid surfaces (such as raindrops on windows) (Dussan V. & Davis, 1974; Dussan V. & Chow, 1983). In technological systems it is harnessed for actuating microfluidic platforms (Atencia & Beebe, 2005; Stone et al., 2004; Whitesides, 2006), for cooling electronic circuits (Amon et al., 2001; Bar-Cohen et al., 2006), for desalination (Fletcher et al., 1974) and for a variety of manufacturing applications (Fendler, 1996; Nagayama, 1996; Wang et al., 2004). Many different mechanisms govern dynamic wetting, where capillary forces at the free surface of a liquid film or a meniscus contribute appreciably to the dynamics of the film. In addition to capillary forces, dynamic wetting may be powered by gravity, thermal or solutal Marangoni effects, or electrokinetics, among others, see Oron et al. (1997); Craster & Matar (2009) for comprehensive reviews. In the present work we focus on the contributions of mechanical vibrations in the solid substrate, particularly MHz-frequency Rayleigh surface acoustic waves (SAWs), to dynamic wetting (Morozov & Manor, 2018). Here, we concentrate on the contribution of Eckart streaming – mass transport that results from the attenuation of sound or ultrasound waves in the bulk of the liquid film. While extensively observed and described in the scientific literature (Shiokawa et al., 1989; Brunet et al., 2007, 2009; Yeo & Friend, 2014), dynamic wetting of a solid substrate by a liquid film, induced by Eckart streaming, was not studied in detail to date. In this work we use thin film (long wave) theory alongside physical experiments to elucidate the phenomenon.
There are three types of acoustic streaming mechanisms by which acoustic waves or waves at the free surface of a liquid or a solid boundary (surface acoustic waves) generate streaming – steady flow along the path of the wave. Firstly, a wave – acoustic wave or surface wave – in the solid or in liquid, invokes periodic viscous flow within the viscous penetration length $\delta\equiv\sqrt{2\mu/\rho\omega}$ away from the interface, where $\mu,~{}\rho$ and $\omega$ are the viscosity and density of the liquid and the wave angular frequency, respectively. Examples in which this problem was considered include the work by Rayleigh on a standing acoustic wave grazing a solid bed; by Schlichting (1932) on standing vibration of waves of infinite and finite (respectively) wavelength in a solid along its surface; by Longuet-Higgins (1953) on shallow ocean waves; and by Manor et al. (2012) on propagating Rayleigh (surface acoustic) waves in a solid substrate. Convective contributions due to surface waves invoke drift of liquid mass, which does not attenuate away from the solid surface and is known variously as the Stokes drift (Stokes, 1847), the Rayleigh law of streaming (Lighthill, 1978), and Schlichting streaming (Yeo & Friend, 2014); although the physics and mathematical modeling of the cases studied by Stokes, Rayleigh, and Schlichting are similar, Rayleigh is the only one among them that explicitly considered acoustic waves.
Secondly, in the presence of an acoustic wave in the liquid, whether due to an actuator or an acoustic wave in the solid, one should expect acoustic radiation pressure (Campbell, 1970; Shiokawa et al., 1989). Convective effects – Reynolds stresses – within the wave increase the local steady pressure in the bulk of the liquid; upon reflection of the wave at an interface between phases of mismatched acoustic impedance (density multiplied by acoustic phase velocity), the wave imposes convective stress, known as the acoustic radiation pressure (King, 1934; Borgnis, 2010; Hasegawa et al., 2000), which is capable of deforming and displacing the interface (Alzuaga et al., 2005; Biwersi et al., 2000; Rajendran et al., 2022).
Finally, a third (and the most well-known) mechanism that may appear due to the attenuation of sound or ultrasound waves in fluid is Eckart streaming (Eckart, 1948; Lighthill, 1978; Nyborg, 1952), where wave intensity is converted to fluid momentum, characterized by an intense vortical flow field that appears in the bulk fluid. This is the main acoustic mechanism for the actuation of fluid in micro-channels (Yeo & Friend, 2014; Wixforth et al., 2004; Wixforth, 2003) and in drop microfluidics (Guttenberg et al., 2004; Brunet et al., 2007, 2009, 2010).
These different acoustofluidic mechanisms govern mass transport at different length scales (Rezk et al., 2014). In the presence of sound or ultrasound waves in the fluid, the surface effect, i.e., Stokes drift, governs the transport of mass at separations from the solid that are smaller than the acoustic wavelength in the fluid. Further away, mass transport in the fluid is governed by Eckart streaming. Acoustic radiation pressure effects may be relevant in the former case, and also in the latter case when Eckart streaming is weak. Moreover, SAW diffraction adds an additional spatial dimension to flow actuation by SAW. While the SAW wavelength usually determines the lower spatial resolution of the resulting acoustic flow, SAW diffraction allows for breaking the SAW wavelength limit as a measure for spatial flow resolution (Fakhfouri et al., 2018).
There are abundant studies on the dynamic wetting of thin liquid films that are governed by Stokes drift and acoustic radiation pressure. Rayleigh surface acoustic waves (SAWs) were recently shown to power the continuous spreading of sub-micrometer to hundreds of micrometer thick films of fully wetting silicon oils (Rezk et al., 2012, 2014) and partially wetting water/surfactant mixtures (Altshuler & Manor, 2015, 2016; Mhatre et al., 2016) along or opposite the path of the propagating SAW; water mixtures evaporate leaving behind SAW-induced deposit morphology. However, we are not aware of studies that consider in detail the acoustic/capillary balance of dynamic wetting in liquid films that are sufficiently thick that mass transport therein, and hence film dynamics, are actuated by Eckart streaming.
In this paper, we use physical experiments, theoretical modeling and simulations to elucidate the physics of dynamic wetting powered by Eckart streaming under the action of MHz-frequency Rayleigh-type SAWs. We describe our experimental findings in Sec. 2 and then proceed with describing the theoretical model in Sec. 3. Theoretical results are presented in Sec. 4; we first consider computational results in Secs. 4.1 and 4.2 and then proceed with an exploration of translationally-invariant solutions in Sec. 4.3. Section 5 is devoted to the summary and conclusions.
2 Experiment
[…]
Figure 1: (A) Upper schematic view of the experimental setup: Surface acoustic wave (SAW) propagates from the interdigitated transducer (IDT) until it reaches the acoustic absorber (comprised of glycerol-soaked paper placed on the actuator under and to the right of the needle). The needle is of known diameter 510 $\mu$m, placed for identifying spatial resolution in the images. (a-e) Successive snapshots taken from an experiment monitoring the flow of a silicone oil film. During the experiment, a drop of silicone oil is placed on the horizontal surface (a), deforms due to the application of SAW (b), and moves in the direction of SAW propagation (c-e).
[…]
Figure 2: Two sets of images from experiments monitoring silicone oil powered by SAWs of three different amplitudes (1.2 nm-1.4 nm), where (A-C) silicone oil starts from the same location, moves in the direction of the SAW, and reaches different distances (a-c) after 4.09 seconds. The corresponding side-view videos at different acoustic power levels are used to analyze the speed and profile of silicone oil during the movement. (D, d) show top view snapshots (corresponding to (C,c)) of the oil. A thin film of oil, hardly visible in the side view, can be seen more clearly behind the main body of silicone oil in the top views.
Figure 3: Successive snapshots taken from the baseline experiments monitoring silicone oil spreading without a SAW, where the outline of the oil drop is marked as a red dashed line. The time/radius stamps are given at the top right corner of each figure part.
Figure 4: (a) Time evolution of the front position for several SAW amplitudes for $A=0.52$ nm (black), $0.65$ nm (gray), $0.78$ nm (purple), $0.91$ nm (cyan), $1.04$ nm (green), $1.17$ nm (magenta), $1.30$ nm (red), $1.43$ nm (brown), $1.56$ nm (orange), $1.69$ nm (blue). The arrow points the direction of increasing $A$. (b) Time variations of the maximum drop height for SAW amplitudes $A=0$ nm (green, no SAW applied), $0.5$ nm (blue), $0.7$ nm (orange), $1.3$ nm (yellow) and $1.6$ nm (purple). The symbols correspond to the measurements and the lines are simple fits to guide the eye. We expect errors of ±$\pm 50~{}\mu$m in the values of the film height due to the limited resolution of the side view image and tracker software.
In our experiments, we vary the SAW displacement amplitude at the surface of the actuator by varying the applied voltage level input to the SAW device (there is a linear relationship between the two; see Appendix A for the corresponding calibration graph). To confirm that the observed dynamics are indeed induced by SAW, we also carried out control experiments, shown in Fig. 3, in which SAW is absent and all other parameters are kept the same. The oil puddle in this experiment, which was repeated three times, slowly spreads concentrically, as expected in the absence of a directional forcing mechanism.
We then employ the side view video, recorded by a goniometer camera, to capture the position of the advancing three-phase contact line of the SAW-actuated oil film using the open-source software Tracker (Brown & Cox, 2009) (see Fig. 4(a) for typical data). We also measure the maximum height of the moving oil film using the public domain image processing software ImageJ (typical results in Fig. 4(b)). Figure 4(a) shows that, as expected, the larger the SAW amplitude $A$, the faster the oil film moves due to the stronger induced leakage wave (Shiokawa et al., 1989). Another notable aspect of SAW-induced dynamics is the contact line speed for early times ($<1$ s): initially, it is rather small, and later it increases and then remains constant. This change of contact line speed appears to be a consequence of the initial change of shape of the drop (see also Fig. 1); while this is taking place, contact line speed is small. After the initial deformation, the drop shape remains unchanged, translating approximately uniformly (this will be seen in Fig. 9 later in the text). Figure 4(b) shows consistent results for the maximum drop height, with an increase of the height for early times, and constant values for longer times. Note that this uniform translation and constant film height differ significantly from the observed behavior for spreading under a body force such as gravity (for gravity-driven spreading such as in Fig. 3 the contact line speed decreases, and the drop becomes thinner as it spreads). These experimental observations will be explained and discussed within the framework of our theoretical model, which is described next.
3 Model
To model the experiments described in Sec. 2, we consider a fluid traversed by a sound wave. The fluid velocity, $\mathbf{V}$, can be written as
$\mathbf{V}=\mathbf{v}+\mathbf{v}_{s},$ (1)
where $\mathbf{v}$ is the “bulk” velocity of the fluid, associated with a characteristic time of the flow, $t_{c}$, which is on the order of seconds (so $|\mathbf{v}|\sim\ell/t_{c}$, if $\ell$ is a typical lengthscale of the flow), and $\mathbf{v}_{s}$ is associated with the much faster timescale of the sound wave, $T=f^{-1}\ll t_{c}$ (typically $T\sim 10^{-6}$ s). The idea behind Eq. (1) is that $\mathbf{v}_{s}$ corresponds to an oscillatory flow generated by the propagating sound wave, while $\mathbf{v}$ is the velocity due to external forces (or to the wave propagation itself, as we will show below). Therefore, the fast-time averages, $\langle\cdot\rangle=(1/T)\int_{0}^{T}\cdot\,dt$, of $\mathbf{v}$ and $\mathbf{v}_{s}$, are given by
$\langle\mathbf{v}\rangle\approx\mathbf{v},\qquad\langle\mathbf{v}_{s}\rangle=0.$ (2)
We write now the Navier–Stokes (NS) equations for the fluid in the compact form
$\frac{\partial(\rho\mathbf{V})}{\partial t}+\rho\left(\mathbf{V}\cdot\nabla\right)\mathbf{V}+\mathbf{V}\,\nabla\cdot\left(\rho\mathbf{V}\right)=-\nabla p+\mu\nabla^{2}\mathbf{V}+\left(\zeta+\frac{\mu}{3}\right)\nabla\left(\nabla\cdot\mathbf{V}\right),$ (3)
where $\rho$ is the density of a compressible fluid and $\mu$, $\zeta$ are the first and second dynamic viscosities, respectively. Since we seek the resulting flow field for large times (of the order of $t_{c}$), we take the fast-time average of Eq. (3) as defined above. Thus, using Eqs. (1) and (2), we find
$\frac{\partial(\rho\mathbf{v})}{\partial t}=-\nabla p+\mathbf{F}_{s}+\mu\nabla^{2}\mathbf{v}+\left(\zeta+\frac{\mu}{3}\right)\nabla\left(\nabla\cdot\mathbf{v}\right),$ (4)
where
$\mathbf{F}_{s}=-\langle\rho_{0}\left(\mathbf{v}_{s}\cdot\nabla\right)\mathbf{v}_{s}+\mathbf{v}_{s}\nabla\cdot\left(\rho_{0}\mathbf{v}_{s}\right)\rangle,$ (5)
and $\rho_{0}=\langle\rho\rangle$ is the average (constant) density of the fluid (note that $\langle\partial(\rho_{0}\mathbf{v}_{s})/\partial t\rangle=0$). Equation (4) shows that the flow field $\mathbf{v}$, which develops on the long time scale $t_{c}$, satisfies NS equations similar to those for $\mathbf{V}$, but with an additional volume force, $\mathbf{F}_{s}$. This flow field, $\mathbf{v}$, is known as acoustic streaming flow, resulting from the high-frequency sound wave traversing the fluid (see, e.g. Shiokawa et al. (1989)).
When the sound wave is a leaky wave generated at a solid substrate that propagates in the $(x,z)$–plane (see Fig. 5), the fluid particle displacement, $\mathbf{u}$, can be written as (Shiokawa et al., 1990)
$u_{x}$ $=$ $\Re\left[A\,e^{\imath(k_{L}x^{\prime}-\omega t)+\imath\alpha_{1}k_{L}z}\right]=A\,e^{-k_{i}x^{\prime}-\alpha_{1}k_{i}z}\cos\left[k_{r}x^{\prime}+\alpha_{1}k_{r}z-\omega t\right],$ (6)
$u_{y}$ $=$ $0,$ (7)
$u_{z}$ $=$ $\alpha_{1}u_{x},$ (8)
where $A$ is the wave amplitude, $k_{L}=k_{r}+\imath\,k_{i}$ is the complex wavenumber of the leaky wave and $x^{\prime}=x-x_{\rm ref}(t)$, where $x_{\rm ref}(t)$ is a reference $x$–coordinate determining the onset of attenuation, whose origin will be discussed in Sec. 4. Also, we define
$\alpha_{1}=\sqrt{(c_{s}/c_{\rm oil})^{2}-1},$ (9)
where $c_{s}$ and $c_{\rm oil}$ are the SAW phase velocities in the substrate and silicon oil, respectively. By choosing (Roditi, 2023) $c_{s}=3880$ m/s and (PrecisionFlow, 2023) $c_{\rm oil}=1350$ m/s, we obtain that $\alpha_{1}=2.69$. Thus, the Rayleigh angle, $\theta$, is given by $\theta=\arctan\alpha_{1}=69.65^{\circ}$. The SAW attenuation factor is given by (Arzt et al. (1967))
$k_{i}=\frac{1}{l_{i}}=\frac{\rho_{\rm oil}}{\rho_{s}}\frac{c_{\rm oil}}{c_{s}^{2}}f,$ (10)
[…]
Figure 5: Schematic of a liquid drop (blue curve) forced by a SAW traveling from left to right, with rear and front contact lines at $x_{r}(t)$ and $x_{f}(t)$, respectively. The thin red line represents the amplitude $A$ of the leaky SAW at the substrate and shows how it is attenuated by liquid (the black dashed line is the envelope of the decaying amplitude in the liquid). The Rayleigh angle is denoted by $\theta$.
The fluid particle displacement in Eqs. (6)–(8) is related to the Eulerian velocity field, $\mathbf{v}_{s}=(v_{s,x},v_{s,y},v_{s,z})$, by
$v_{s,x}=\frac{du_{x}}{dt},\qquad v_{s,y}=\frac{du_{y}}{dt},\qquad v_{s,z}=\frac{du_{z}}{dt},$ (11)
which, on substitution into Eq. (5), leads to explicit expressions for the force components,
${\rm F}_{s,x}$ $=$ $\left(1+\alpha_{1}^{2}\right)k_{i}\rho_{0}A^{2}\omega^{2}e^{-2k_{i}\left(x^{\prime}+\alpha_{1}z\right)},$ (12)
${\rm F}_{s,y}$ $=$ $0,$ (13)
${\rm F}_{s,z}$ $=$ $\alpha_{1}F_{s,x}.$ (14)
The force $\mathbf{F}_{s}$ is conservative, i.e., it can be expressed as $\mathbf{F}_{s}=-\nabla\phi_{s}$ where the potential, $\phi_{s}$, is given by
$\phi_{s}(x,z)=\phi_{0}\,e^{-2k_{i}\left(x^{\prime}+\alpha_{1}z\right)},\quad\mbox{with}\quad\phi_{0}=\frac{1+\alpha_{1}^{2}}{2}\rho_{0}A^{2}\omega^{2}.$ (15)
Therefore, the NS equations for the velocity field $\mathbf{v}$, under the influence of both SAW and gravitational forces, can be written as
$\rho_{0}\frac{\partial\mathbf{v}}{\partial t}=-\nabla P+\mu\nabla^{2}\mathbf{v},$ (16)
where we assume that the $\mathbf{v}$ field is incompressible ($\nabla\cdot\mathbf{v}=0$), and we have defined the effective pressure, $P$, as
$P=p+\phi_{s}+\phi_{g},$ (17)
with $\phi_{s}$ given by Eq. (15) and $\phi_{g}(z)=\rho_{0}gz$, corresponding to a gravitational body force, $\mathbf{F}_{g}=-\nabla\phi_{g}=-\rho_{0}g\,\hat{\mathbf{k}}$, where $\hat{\mathbf{k}}$ is a unit vector in the out-of-plane direction.
In this work, we consider incompressible, low Reynolds number free surface flows under the long-wave approximation so that all inertial terms on the left-hand side of Eqs. (16) are negligible with respect to the viscous and pressure terms. Carrying out a standard long-wave expansion of (16) and imposing the usual stress boundary conditions at the free surface (vanishing tangential stress and normal stress balanced by surface tension, see e.g. Kondic (2003) for details), we obtain the following fourth-order differential equation for the film height, $z=h(x,y,t)$,
$3\mu\frac{\partial h}{\partial t}-\nabla\cdot\left(h^{3}\nabla P\right)=0,$ (18)
where
$P(x,y,t)=-\gamma\nabla^{2}h+\rho_{0}gh+\phi_{s}(x,h).$ (19)
Here, $\gamma$ is the coefficient of surface tension at the oil-air interface, assumed constant.
We nondimensionalize the problem using an arbitrary length scale, $\ell$, a timescale $t_{c}$ such that leading order terms balance in (18), and a pressure scale $p_{c}$ based on the capillary contribution in (19), so that we have
$(x,y)=\ell(\tilde{x},\tilde{y}),\quad(x_{\rm ref},h)=\ell(\tilde{x}_{\rm ref},\tilde{h}),\quad t=t_{c}\tilde{t}=\frac{3\mu\ell}{\gamma}\tilde{t},\quad P=p_{c}\tilde{P}=\frac{\gamma}{\ell}\tilde{P},$ (20)
where the tilded quantities are dimensionless. Thus, the dimensionless governing equation is
$\frac{\partial\tilde{h}}{\partial\tilde{t}}-\tilde{\nabla}\cdot\left(\tilde{h}^{3}\tilde{\nabla}\tilde{P}\right)=0,$ (21)
where
$\tilde{P}=-\tilde{\nabla}^{2}\tilde{h}+{\rm Bo}\,\tilde{h}+{\cal S}\,\tilde{\phi}_{s}(\tilde{x},\tilde{h}),$ (22)
with
$\tilde{\phi}_{s}(\tilde{x},\tilde{h})=e^{-2\tilde{k}_{i}[(\tilde{x}-\tilde{x}_{\rm ref})+\alpha_{1}\,\tilde{h}]},$ (23)
and we have defined $\tilde{k}_{i}=\ell k_{i}$. The dimensionless parameters introduced here are given by
${\rm Bo}=\frac{\rho_{0}\ell^{2}}{\gamma}g=\frac{\ell^{2}}{a^{2}},\qquad{\cal S}=\frac{\ell}{\gamma}\phi_{0},$ (24)
where $a=\sqrt{\gamma/(\rho_{0}g)}$ is the capillary length and $\phi_{0}$ was defined in (15). In our simulations, motivated by a typical drop size in the experiments, we set $\ell=1$ mm. For simplicity, we omit the tilde from now on so that all variables will be dimensionless unless specific units are explicitly written, mainly when comparison with experimental data is required.
4 Theoretical results
In this section, we discuss the predictions of the model developed in Sec. 3, and compare them with the experimental results. In Sec. 4.1 we present results for a chosen reference case, and then in Sec. 4.2, we discuss the influence of model parameters, in particular the acoustic amplitude, on the results. In Sec. 4.3 we discuss (approximate) traveling-wave solutions. All simulations are carried out in two physical dimensions, assuming translational invariance in the transverse $y$-direction. The discussion of the influence of the three-dimensional configuration will be presented elsewhere.
4.1 Main features of the results
In our simulations, we solve Eq. (21) numerically using COMSOL™; see Appendix B for the description of the implementation. As the initial condition, we consider the two–dimensional parabolic drop of cross-sectional area
[…]
which remains constant during the drop evolution. The drop height parameter $h_{d}$ is obtained from the known experimental (3D) drop volume, $V_{d}$, as $h_{d}=2V_{d}/(\pi r_{d}^{2})$, using the formula for the volume $V_{d}$ of a 3D parabolic cap of base radius $r_{d}$ and height $h_{d}$. The 2D parabolic initial condition is centered at $x=x_{d}$, and is written as
$h(x,0)=\left\{\begin{array}[]{ll}h_{p},&x<x_{r}(0),\\ \left(h_{d}-h_{p}\right)\left[1-\left(\frac{x-x_{d}}{r_{d}}\right)^{2}\right]+h_{p},&x_{r}(0)\leq x\leq x_{f}(0),\\ h_{p},&x>x_{f}(0),\end{array}\right.$ (26)
where $x_{r}(0)=x_{d}-r_{d}$ and $x_{f}(0)=x_{d}+r_{d}$ are the initial positions of the rear and front contact lines of the drop, respectively (see Fig. 5). Here, $h_{p}$ ($\ll h_{d}$) is the thickness of a precursor film, introduced for numerical reasons in order to avoid the stress singularity associated with a moving contact line. It is well known (see, e.g. Diez & Kondic (2002)), that the step size $\Delta x$ in the spatial discretization should be similar to $h_{p}$ to ensure numerical convergence, and therefore we use $\Delta x=h_{p}$ in our simulations. The specific value assigned to $h_{p}$ has only a minor influence on the results; e.g., if $h_{p}$ is doubled or halved, the change of spreading speed or drop height is just 1-2%. All reported results are obtained using $h_{p}=0.01$.
Recalling the introduction of the reference $x$-coordinate origin, $x_{\rm ref}(t)$, in Sec. 3, we note that the SAW experiences minimal attenuation in regions of the actuator where fluid is not present, with significant attenuation only under the bulk droplet. To model this, we assume that the acoustic potential, $\phi_{s}$, is constant (no attenuation) in regions where the film height is smaller than some critical thickness $h^{*}$, and attenuates where $h>h^{*}$,
$\phi_{s}=\left\{\begin{array}[]{ll}1,&x<x_{1}^{\ast}(t)\,,\\ e^{-2k_{i}\left[x-x_{1}^{\ast}(t)+\alpha_{1}\left(h-h^{\ast}\right)\right]},&x_{1}^{\ast}(t)\leq x\leq x_{2}^{\ast}(t)\,,\\ e^{-2k_{i}\left[x_{2}^{\ast}(t)-x_{1}^{\ast}(t)\right]},&x>x_{2}^{\ast}(t)\,,\end{array}\right.$ (27)
where $x_{1}^{\ast}$, $x_{2}^{\ast}$ are defined by $h(x_{1}^{\ast},t)=h(x_{2}^{\ast},t)=h^{\ast}$ (this form of $\phi_{s}$ corresponds to the choice $x_{\rm ref}(t)=x_{1}^{*}(t)$). This model is further motivated by the known fact that the Eckart streaming—the acoustic streaming that pushes the oil film along the path of SAW in this work—governs the flow when the oil film is sufficiently thick. Close to the solid surface the flow is governed by a boundary layer flow (known as the Schlichting streaming to some and Rayleigh streaming to others), generated by the SAW in the solid. These mechanisms are observed to support the spreading of $\lambda_{\rm oil}/4$ thick oil films opposite the path of the SAW. However, thicker oil films are governed by Eckart streaming and spread along the path of the SAW. Hence, we account for Eckart streaming-induced acoustic forcing where the film thickness is greater than $h^{\ast}=\lambda_{\rm oil}/(4\ell)$  (Rezk et al., 2012, 2014; Altshuler & Manor, 2015, 2016).
[…]
Experiment Simulation
[…]
$\mu=0.5$ Poise ${\rm Bo}=0.452$
$\gamma=20.8$ dyn/cm $A_{d}=2.12$
$\omega=2\pi\times 20$ MHz $r_{d}=3.2$
$A=1.3$ nm $h_{d}=0.497$
$l_{i}=2.7$ mm $k_{i}=0.37$
[…]
Figure 6(a) shows the evolution of the thickness profile obtained by solving Eq. (21) for the parameter values in Table 1, subject to the initial condition given by Eq. (26), and boundary conditions fixing the film thickness to $h_{p}$ and requiring vanishing derivatives at the domain boundaries. We observe that the simulated drop settles to a steady shape, which translates uniformly. The steady shape of the bulk droplet has a pronounced “foot” at its rear; though we do not yet fully understand the details of why this forms, it is consistent with the overall drop asymmetry that develops, seen in Fig. 2. We discuss these features of the results further below.
Figure 6(b) shows a single thickness profile at $t=2$ s (blue curve), which is compared with the experimental data (black symbols). We note that capturing the spreading experimental drop geometry (and velocity) requires image analysis of the captured image data; the analysis adds uncertainty to the measured data estimated to be ±$\pm 50$ $\mu$m. Moreover, repeating experiments show some sensitivity to the initial condition, i.e., the thickness of the oil drop placed upon the SAW device at the time SAW is initiated, albeit in all cases the oil reaches a constant speed at long times.
Figure 6: (a) Thickness profiles from $t=0$ (red line) to $t=6$ s every $\Delta t=0.5$ s for SAW amplitude $A=1.30\,$nm. Here, we use $h^{\ast}=17\mu{\rm m}/h_{c}\approx(\lambda_{\rm oil}/4)/h_{c}$. (b) Comparison between simulated (solid blue line) and experimental (black symbols) thickness profiles at $t=2$ s.
4.2 Parametric dependence
In this section, we separate the discussion into two parts: first, we consider early-time behavior, and then long-time evolution. Figure 7 shows the computational results for the maximum drop height $h_{\rm max}$ for different values of the acoustic amplitude, $A$. We show the results for two different initial conditions, with the purpose of illustrating the relevance of the initial drop shape. Focusing first on the initial condition used for Fig. 6 (solid curves in Fig. 7) with initial drop height of $h_{d}\approx 0.5$ mm obtained from the experiments (see also Eq. (25)), we note different behavior for small and large values of $A$, as could be anticipated from the experimental results of Fig. 4(b). For the larger $A$-values, $h_{\rm max}$ initially increases rapidly but then decreases. For the two smallest $A$-values, this initial rapid increase is not observed. In all cases, for longer times, $h_{\rm max}$ slowly decreases towards a constant ($A$-dependent) value, although it does not completely saturate for the times shown. One interpretation of these results is that for large $A$, the acoustic forcing dominates capillary stresses and gravity, leading to an increase of the drop height, while for smaller $A$, the acoustic forcing is not strong enough to lead to initial drop height increase. For the other initial condition (smaller initial drop height, shown by dashed lines in Fig. 7), the details are different, but we again observe an initial increase of $h_{\rm max}$ for sufficiently large values of $A$, followed by (in this case) a slight decrease towards constant $A$-dependent values at large times.
Figure 7: Maximum fluid thickness, $h_{\rm max}$, as a function of time at different SAW amplitudes, $A$, for two initial conditions of the drop: $r_{d}=3.2$ mm (solid lines) and $r_{d}=3.9$ mm (dashed lines). The cross–sectional fluid area, $A_{d}$, is the same, but the latter has a lower initial maximum thickness than the former (see Eq. (25)). The values of are $A$ are: $2.00$ nm (black), $1.69$ nm (blue), $1.30$ nm (red), $1.04$ nm (magenta), $0.65$ nm (brown), $0.20$ nm (green) (same colors correspond to same values of $A$).
Next, we shift our attention to long-time behavior, focusing first on the dependence of drop profiles on SAW amplitude $A$. Figure 8 shows the long-time drop profiles together with the pressure specified by Eq. (22) for four values of $A$. While the maximum drop height increases with $A$, as expected, we note that the thickness of the ‘foot’ directly behind the front increases with $A$ as well (varying between approximately 50 and 100 $\mu$m for the considered values of $A$). In addition, we note a very thin (on the order of the precursor scale, but a little larger) trailing film left behind the translating drop. While both the ‘foot’ and the thin trailing film deposited behind the drop still need to be fully analyzed and understood, we suspect that these features of the results are related to the pressure (plotted by the red lines with values shown on the right-hand axis in Fig. 8); note that the ‘foot’ forms in the region where the pressure is O(1), shown by the dashed lines. On a different note, we point out the strong decrease of the pressure due to the SAW attenuation, as the film height increases from the precursor scale, as well as the pressure dip at the front contact line, caused by the rapidly changing curvature there.
(a) $\,A=1.04$ nm
(b) $\,A=1.30$ nm
(c) $\,A=1.69$ nm
(d) $\,A=2.00$ nm
Figure 8: Simulated drop profiles (blue curves, left axis) and dimensionless pressure $\tilde{P}=P/p_{c}$ (red curves, right axis) at $t=6$ s for four different values of SAW amplitude $A$. The horizontal dashed line is at $\tilde{P}=1$.
Focusing next on the spreading speed, Fig. 9 illustrates that for long times both experiments and simulations show approximately constant spreading speed (drop front position $x_{f}$ increasing linearly in time), which seems to be a characteristic feature of the considered acoustic driving (note that we shift the experimental results in the figure to illustrate better the long time behavior; initial transients may account for differences at early times due to the particulars of the initial conditions). These results suggest that the spreading drop essentially translates without much of shape change for late times. Figure 10(a) confirms this expectation: this figure shows that the width of the spreading drop, ${\rm w}=x_{2}^{*}-x_{1}^{*}$, reaches a plateau and remains constant in the case of larger acoustic amplitudes $A$; on sufficiently long time scales (not shown in the figure), the $\rm w$-values obtained for smaller $A$-values reach a plateau as well.
Figure 9: Comparison between numerical (solid lines) and experimental (symbols) results for the SAW amplitude $A=1.69\,$nm (blue), 1.301.301.30\,1.30nm (red) and 1.041.041.04\,1.04nm (magenta). We have added arbitrary $x$–shifts to the simulation results to facilitate the comparison of the slope of $x_{f}$ versus $t$ between experiments and simulations for long times.
Figure 10: (a) Numerically calculated drop width, $\rm w$, as a function of time for several values of $A$: $=0.20$ nm (green), $=0.65$ nm (brown), $=1.04$ nm (magenta), $=1.30$ nm (red), $=1.69$ nm (blue), and $=2.00$ nm (black). Note that the two smallest values of $A$ have not reached an asymptotic value of $\rm w$ for $t=6$ s, but will do for later times. (b) Long time values of the front speed, ${\rm v}_{f}$ (in mm/s, blue filled circles), maximum thickness, $h_{\rm max}$ (mm, red hollow squares), and drop width, $\rm w$ (cm, black filled squares), as functions of $A$. The points correspond to the raw data, and the solid lines connecting them guide the eye.
To summarize our numerical results, Fig. 10(b) plots the numerically-predicted long-time values (reached for $t\gtrsim 6$ s) of the front velocity, ${\rm v}_{f}$, the maximum drop height, $h_{\rm max}$, and the drop width, ${\rm w}=x_{2}^{\ast}-x_{1}^{\ast}$, as a function of $A$. We note that, while ${\rm v}_{f}$ strongly increases with $A$, the dependence of $h_{\rm max}$ and $\rm w$ on $A$ is much weaker. These results suggest that a good qualitative description of the problem might be obtained by considering a traveling wave type of solution. Such an approach is discussed in the following section.
4.3 Traveling wave solution
In view of the results presented so far, we now seek a traveling wave solution of Eq. (21). Within this framework, we assume that the whole drop translates by constant speed $U$, so that
$x_{r}(t)=Ut,\qquad x_{f}(t)={\rm w}+Ut,$ (28)
where (constant) $\rm w$ is the width of the moving drop. Here, we will consider that the rear (front) contact lines of the traveling drop are given by $x_{1}^{\ast}$ ($x_{2}^{\ast}$) (see Fig. 5), since the SAW force that drives it corresponds to $h>h^{\ast}$. Then, we define $\xi=x-x_{1}^{\ast}(t)=x-Ut$ and assume that $h(x,t)=h(\xi)$, so that Eq. (21) becomes
$U\frac{dh}{d\xi}+\frac{d}{d\xi}\left[h^{3}\frac{d}{d\xi}\left(-\frac{d^{2}h}{d{\xi}^{2}}+{\rm Bo}\,h+S\phi_{s}(\xi,h)\right)\right]=0$ (29)
where
$\phi_{s}(\xi,h)=e^{-2k_{i}[\xi+\alpha_{1}(h-h^{\ast})]}.$ (30)
This equation can be integrated once to yield
$Uh+h^{3}\frac{d}{d\xi}\left[-\frac{d^{2}h}{d{\xi}^{2}}+{\rm Bo}\,h+S\phi_{s}(\xi,h)\right]=J,$ (31)
where $J$ represents the flux. The traveling wave solution must be calculated for $0\leq\xi\leq\xi_{f}$, along with the following boundary conditions at $\xi=0$
$h(0)=h^{\ast},\qquad h^{\prime}(0)=h^{\prime\prime\prime}(0)=0,$ (32)
where the prime denotes $d/d\xi$. The values of $h^{\prime\prime}(0)$, $U$ and $\xi_{f}$ are determined by
[…]
Note that $J$ corresponds to the flux within a thin film of thickness $h^{\ast}$ that enters the drop at $\xi_{f}$ and comes out of it at $\xi=0$ (with respect to the reference frame fixed at the drop). Also, note that $\xi_{f}$ is exactly the width of the traveling drop, $\rm w$.
We develop an iterative scheme to perform the numerical integration of Eq. (31), under the conditions given by Eq. (33). We start by guessing values of $h^{\prime\prime}(0)$, $U$ and $\xi_{f}$ to perform the integration of Eq. (31) and modify them accordingly until these conditions are satisfied within a small relative error (typically, $10^{-7}$). After obtaining convergence for a given $A$, we use these converged values of $h^{\prime\prime}(0)$, $U$ and $\xi_{f}$ as new guess values for $A+\Delta A$ (continuation). $\Delta A$ has to be very small to get a new convergence (typically, $\Delta A=10^{-5}\,$nm). Figure 11(a) shows the results of drop profiles for the values of $A$ considered so far. Clearly, drops became taller and narrower as $A$ increases. To compare the results with the experimental ones, we note that for $A=1.3$ nm, the dimensional drop speed is ${\rm v}_{f}=(\ell/t_{c})U=2.75$ mm/s, which is of the order of the experimental value. As expected, the value of $J$ is small compared to unity, $J=3.26\times 10^{-4}$. Figure 11(b) shows the dependence of drop speed, ${\rm v}_{f}$, its width, $\rm w$, and the maximum thickness, $h_{\rm max}$, on the SAW amplitude, $A$. Note that the model predicts that the speed depends approximately quadratically on the amplitude, $A$, as expected since the drop thickness increases linearly with $A$. We note that the trends shown in this figure for the travelling wave are similar to those in Fig. 10(b) obtained using full numerical solution.
This approximate formulation shows that the simple traveling-wave solution can capture, at least qualitatively, features of the experimental drops. Clearly, however, this description cannot account for the rear flat film left behind the drop and the fact that the actual rear contact line remains practically at rest.
Figure 11: (a) Thickness profile of the traveling wave solution for: $A=0.50$ nm (green), $0.65$ nm (black), $1.04$ nm (magenta), $1.30$ nm (red), $1.69$ nm (blue). (b) Front speed, ${\rm v}_{f}$ (mm/s, black line), drop width, $\rm w$ (mm, blue line), and maximum height, $h_{\rm max}$ (mm, red line), of the traveling drop as a function of the SAW amplitude, $A$.
We note that we can use these results also to obtain an approximate value of $U$ by applying a simple formula to estimate the scale of the front speed, ${\rm v}_{f}$. By setting $J\approx 0$ and neglecting both capillary and gravity effects in Eq. (29), we have
$U\approx-Sh^{2}\frac{d\phi_{s}}{d\xi}=2k_{i}Sh^{2}e^{-2k_{i}[\xi+\alpha_{1}(h-h^{\ast})]}.$ (34)
By setting $h=h_{d}/\ell$ and $\xi=2r_{d}/\ell$, this expression yields $v_{f}=U\ell/t_{c}=0.4$ mm/s for $A=1.3$ nm, which is the right order of magnitude.
5 Conclusions
This paper presents experimental results and a self-consistent model describing the spreading of a silicone oil drop under the influence of acoustic driving produced by a surface acoustic wave (SAW). Both simulations and the theoretical model show consistent features, and in particular constant spreading speed for long times. This feature of the results is specific to this type of forcing, and is different from spreading under a body force (such as gravity) where fluid thickness slowly decreases as a drop spreads, leading to decreased spreading speed (which typically scales as $h^{2}$, where $h$ is the drop thickness). Our understanding is that this feature of the results is a consequence of the particular forcing imposed by SAW, which decreases as a function of film height; therefore, in contrast to the gravity spreading, here we have a situation where the drop is essentially pushed from behind, leaving just a very thin (on the order of the precursor scale) film behind the spreading drop. The drop profile itself includes a ‘foot’ behind the main drop; this feature of the results still needs further investigation.
While not all details of the computational and experimental results are fully understood, in general, we find very reasonable agreement between the experiments and theoretical results, in particular regarding drop shapes, spreading speeds, and the dependence of the results on the SAW amplitude. For long times, both experiments and simulations show that a drop essentially translates with constant speed and without change of shape. This finding motivated us to formulate a traveling wave description of the system. The resulting solution predicts a linear increase in the drop thickness and a quadratic increase of the spreading speed with the acoustic amplitude $A$.
While many features of the experimental results have been rationalized using the formulated model, we note that there is still much to be done. In particular, further investigation is required to understand all the features of the calculated drop profiles, and also three-dimensional aspects of the results remain to be simulated and understood. Consideration of such effects will be the subject of future work.
Appendix A Calibration of applied voltage versus SAW displacement
We use the normal displacement amplitude (particle displacement of the SAW), $A$, to represent the SAW strength. In the experiments, we control this amplitude by changing the applied voltage of the signal generator. Therefore, we conducted an experiment to obtain a calibration curve between the applied voltage and amplitude, as shown in Fig 12. We measure $A$ over a surface of $1\times 1~{}\text{mm}^{2}$ about 2 mm away from the SAW actuator using laser doppler vibrometer. We find an approximately linear dependence of $A$ on voltage, consistently with the literature (Ballantine Jr et al., 1996).
Figure 12: A calibration plot of corresponding SAW displacement when different voltages are applied to the actuator.
Appendix B Computational implementation
Equation (21) is written in a form convenient for the use of COMSOL™Multiphysics PDE Coefficients Form. This package solves, by finite elements, a vectorial equation for the unknown vector $\vec{u}=(u_{1},u_{2},\dots,u_{N})^{\top}$. The equation is of the form
${\bf e}\frac{\partial^{2}\vec{u}}{\partial t^{2}}+{\bf d}\frac{\partial\vec{u}}{\partial t}+\nabla\cdot\left(-{\bf c}\nabla\vec{u}-{\bf\alpha}\,\vec{u}+{\bf\gamma}\right)+{\bf\beta}\,\nabla\vec{u}+{\bf a}\,\vec{u}=\vec{f},$ (35)
where the coefficients of the $N$ scalar equations are in the matrices ${\bf e}$, ${\bf d}$, ${\bf\gamma}$, ${\bf a}$ (of dimensions $N\times N$), ${\bf\alpha}$, ${\bf\beta}$ (of dimensions $N\times N\times n$), ${\bf c}$ (of dimensions $N\times N\times n\times n$) and the vector $\vec{f}$ (of dimension $N$), where $n$ is the spatial dimension of the problem ($n=1,2,3$). In index notation, this equation reads as
$e_{ij}\frac{\partial^{2}u_{j}}{\partial t^{2}}+d_{ij}\frac{\partial u_{j}}{\partial t}+\frac{\partial}{\partial x_{l}}\left(-c_{ijkl}\frac{\partial u_{j}}{\partial x_{k}}-\alpha_{ijl}u_{j}+\gamma_{il}\right)+\beta_{ijl}\frac{\partial u_{j}}{\partial x_{l}}+a_{ij}u_{j}=f_{i},$ (36)
where $i,j=1.\ldots,N$ and $k,l=1,\ldots,n$.
The considered system, given by Eq. (21), contains two components $u=(\tilde{h},\tilde{P})$ ($N=2$) and we use two equations, namely, Eqs. (21) and (22) for ($n=1)$, corresponding to the solution depending on a single spatial variable, $x$.
In the following, we list the non-vanishing coefficients (we omit the indexes $k$ and $l$ for brevity and consider $x_{1}\equiv x$ since $k=l=1$):
  • Row $1$ ($i=1$) for Eq. (21)
    $d_{11}=1,\quad c_{12}=\tilde{h}^{3}.$ (37)
  • Row $2$ ($i=2$) for Eq. (22)
    $c_{21}=-1,\quad a_{21}=-{\rm Bo},\qquad a_{22}=1,\qquad f_{2}={\cal S}{\tilde{\phi}}_{s}$ (38)
At the domain ends, we apply Dirichlet boundary conditions $\tilde{h}=\tilde{h}_{p}$, where $\tilde{h}_{p}$ is the precursor film thickness, and $\partial h/\partial x=0$. In order to achieve convergent numerical simulations, we discretize the domain with steps $\Delta x=h_{p}$.
Declaration of Interests
The author report no conflict of interest.